OpenStudy (josedavid):

The endpoints of segment EF are E(-2, 3) and F(5, -3). Find the coordinates of the midpoint of EF. (1 ½, 0) (½, 2) (0, 5) (3, 0)

OpenStudy (agent0smith):

Use the midpoint formula (google if you don't know it), plug in the values of x and y.

OpenStudy (josedavid):

IS C OR D

OpenStudy (maebaby7):

Its C:)

OpenStudy (eliesaab):

E(-2, 3) and F(5, -3) \[ \frac{5-2}{2}=\frac 32=1 \frac 12\]

OpenStudy (eliesaab):

\[ \frac{3-3}2=0 \]

OpenStudy (eliesaab):

The answer is A